# Surgical Workflow Recognition and Blocking Effectiveness Detection in Laparoscopic Liver Resections with Pringle Maneuver

### v1

@@ Abstract
Pringle maneuver (PM) in laparoscopic liver resections aims to reduce blood loss and provide a clear surgical view by intermittently blocking blood inflow of the liver, whereas prolonged PM may cause ischemic injury. To comprehensively monitor this surgical procedure and provide timely warnings of ineffective and prolonged blocking, we suggest two complementary AI-assisted surgical monitoring tasks: workflow recognition and blocking effectiveness detection in liver resections. The former presents challenges in real-time capturing of short-term PM, while the latter involves the intraoperative discrimination of long-term liver ischemia states. To address these challenges, we meticulously collect a novel dataset, called PmLR50, consisting of 25,037 video frames covering various surgical phases from 50 laparoscopic liver resection procedures. Additionally, we develop an online baseline for PmLR50, termed PmNet. This model embraces Masked Temporal Encoding (MTE) and Compressed Sequence Modeling (CSM) for efficient short-term and long-term temporal information modeling, and embeds Contrastive Prototype Separation (CPS) to enhance action discrimination between similar intraoperative operations. Experimental results demonstrate that PmNet outperforms existing state-of-the-art surgical workflow recognition methods on the PmLR50 benchmark. Our research offers potential clinical applications for the laparoscopic liver surgery community. Source code and data will be publicly available.
@@ Introduction
To address the above challenges, we propose a unified online baseline called PmNet for both PM workflow recognition and blocking effectiveness detection. Our framework includes a Masked Temporal Encoding (MTE) that emphasizes surgical action details by adaptively masking and swapping message tokens for efficient short-term temporal information modeling. Moreover, we introduce a Compressed Sequence Modeling (CSM) operation, which leverages the state space model (SSM) to create long-range dependencies on temporally pooled long-term features and information of tissue ischemia regions for compressed feature modeling, and then interacts short-term memory with long-term memory to facilitate contextual retrieval. Additionally, our model embeds a Contrastive Prototype Separation (CPS) strategy to expand the feature space distance across target prototypes, improving the discrimination between similar intraoperative actions. We comprehensively evaluate multiple mainstream surgical workflow recognition methods to construct the PmLR50 benchmark.
Experimental results demonstrate that PmNet outperforms previous state-of-the-art methods on the PmLR50 test set, proving potential clinical significance in surgical intervention and assistance. Our main contributions are summarized as:
@@
  • Blocking Effectiveness. Beyond PM workflow recognition, surgeons also provide corresponding binary labels for video frames associated with blocking actions to determine the effectiveness. The ineffective blocking is defined as “inability to successfully intermittently block blood inflow to the liver by knotting the catheter” (Sidaway 2024). In the PmLR50 dataset, we provide a total of 6 surgical cases of ineffective blocking, representing 12% of the total cases.
@@
Dataset Characteristics and Statistics: As depicted in the left of Fig. 3, the majority of the sampled video frames in PmLR50 range between 400 and 600 frames. Among a total of five surgical phases, ‘Resecting’ contains the highest number of frames, accounting for 66.61% of the total frames, significantly higher than the 9.38% for ‘Knotting’ and 5.86% for ‘Releasing’ (see the right of Fig. 3). Notably, our dual-rate frame sampling (i.e., 3 fps and 0.33 fps) ensures more than 30 frames of blocking-related operations per surgical procedure, thus providing more balanced samples for various surgical phases. Further details of PmLR50 are available in the supplementary material.
@@
PM Workflow Recognition Precision $\uparrow$ 73.67 74.45 68.05 67.54 88.99
Recall $\uparrow$ 64.94 61.87 57.36 60.20 77.23
Accuracy $\uparrow$ 93.06 92.96 92.49 92.84 95.89
Jaccard $\uparrow$ 61.35 61.66 56.67 57.98 71.39
@@
Params $\downarrow$ 10.96M 28.76M 24.61M 121.26M 91.77M
Inference Speed (FPS) $\uparrow$ - - - 14.84 26.08
@@
Implementation Details: Our model is trained with the Adam optimizer on two NVIDIA 4090 GPUs. The batch size is set to 8, and the initial learning rate is 3e-5 with a weight decay of 0.0005. We adopt EfficientNet-B3 (Tan and Le 2021) pre-trained on ImageNet (Deng et al. 2009) as the backbone. The parameters of other components in PmNet are randomly initialized. In our experiments, 20 previous consecutive frames with resolution 256$\times$256 are selected as model inputs in an online manner, with the clip window size $w$ set to 4. During training, we apply color jitter and random horizontal flip for data augmentation. To trade off the accuracy and efficiency of PmNet, we embed four self-attention layers and three token swap operations in MTE, with the number of SSM blocks set to 2 in CSM.
@@
Methods Metrics Preparing Knotting Resecting Releasing Postprocess
TeCNO Accuracy 92.62 96.16 87.56 95.33 93.64
Jaccard 68.57 63.66 82.48 64.63 27.41
TMRNet Accuracy 91.64 97.33 85.97 95.91 93.93
Jaccard 62.66 74.46 80.25 68.34 22.60
Trans-SVNet Accuracy 92.08 94.66 89.29 93.72 92.72
Jaccard 65.90 54.45 81.39 54.91 26.74
Surgformer Accuracy 93.96 94.90 89.74 94.18 91.44
Jaccard 69.10 55.84 83.73 56.89 24.33
PmNet Accuracy 97.52 99.02 90.58 98.01 94.29
Jaccard 76.26 89.92 87.34 74.15 29.30
Table 3: Comparison with the state-of-the-arts for PM workflow recognition of each class on the PmLR50 test set.
@@
As illustrated in Table 2, our model shows superior performance in both tasks. Compared to the second-best model TMRNet, our model achieves an improvement of 9.73% Jaccard in PM workflow recognition. Besides, by modeling long-term memory through CSM, PmNet obtains a high accuracy of 98.69% in blocking effectiveness detection. In addition, our model offers certain advantages in terms of parameters and inference efficiency, reaching an inference speed of 26.08 fps for real-time clinical applications. Furthermore, we provide a fine-grained evaluation of the recognition results for each phase in Table 3. PmNet also outperforms other methods across all phases, e.g.Knotting and Releasing, a 15.46% and 5.81% improvement in Jaccard, which can be attributed to CPS to discriminate between confused intraoperative actions. Fig. 5 visualizes the predictions of our method compared to the latest Surgformer for PM workflow recognition, where PmNet demonstrates better consistency and accuracy across all surgical phases.
@@
Ablations for Each Component. Table 4 ablates the contribution of each component of PmNet to surgical workflow recognition. The experimental results indicate that all components, including MTE, CSM, and CPS play a positive role. The results in the last two rows illustrate that MTE used for capturing short-term temporal information shows the most significant effect, increasing the Jaccard value for PM workflow recognition by 12.26%. Meanwhile, the efficient long-term temporal modeling by CSM and the comparative learning of intraoperative activities by CPS also enhance the validity and robustness of PmNet.

### v4

@@ Abstract
Pringle maneuver (PM) in laparoscopic liver resections aims to reduce blood loss and provide a clear surgical view by intermittently blocking blood inflow of the liver, whereas prolonged PM may cause ischemic injury. To comprehensively monitor this surgical procedure and provide timely warnings of ineffective and prolonged blocking, we suggest two complementary AI-assisted surgical monitoring tasks: workflow recognition and blocking effectiveness detection in liver resections. The former presents challenges in real-time capturing of short-term PM, while the latter involves the intraoperative discrimination of long-term liver ischemia states. To address these challenges, we meticulously collect a novel dataset, called PmLR50, consisting of 25,037 video frames covering various surgical phases from 50 laparoscopic liver resection procedures. Additionally, we develop an online baseline for PmLR50, termed PmNet. This model embraces Masked Temporal Encoding (MTE) and Compressed Sequence Modeling (CSM) for efficient short and long-term temporal information modeling, and embeds Contrastive Prototype Separation (CPS) to enhance action discrimination between similar intraoperative operations. Experimental results demonstrate that PmNet outperforms existing state-of-the-art surgical workflow recognition methods on the PmLR50 benchmark. Our research offers potential clinical applications for the laparoscopic liver surgery community.
Codehttps://github.com/RascalGdd/PmNet
@@ Introduction
To address the above challenges, we propose a unified online baseline called PmNet for both PM workflow recognition and blocking effectiveness detection. Our framework includes a Masked Temporal Encoding (MTE) that emphasizes surgical action details by adaptively masking and swapping message tokens for efficient short-term temporal information modeling. Moreover, we introduce a Compressed Sequence Modeling (CSM) operation, which leverages the state space model (SSM) to create long-range dependencies on temporally pooled long-term features and information of tissue ischemia regions for compressed feature modeling, and then interacts short-term memory with long-term memory to facilitate contextual retrieval. Additionally, our model embeds a Contrastive Prototype Separation (CPS) strategy to expand the feature space distance across target prototypes, improving the discrimination between similar intraoperative actions. We comprehensively evaluate multiple mainstream surgical workflow recognition methods to construct the PmLR50 benchmark. Experimental results demonstrate that PmNet outperforms previous state-of-the-art methods on the PmLR50 test set, proving potential clinical significance in surgical intervention and assistance. Our main contributions are summarized as:
@@
  • Blocking Effectiveness. Beyond PM workflow recognition, surgeons also provide corresponding binary labels for video frames associated with blocking actions to determine the effectiveness. The ineffective blocking is defined as “inability to successfully intermittently block blood inflow to the liver by knotting the catheter” (Sidaway 2024). In the PmLR50 dataset, we provide a total of 5 surgical cases of ineffective blocking, representing 10% of the total cases.
@@
Dataset Characteristics and Statistics: As depicted in the left of Fig. 3, the majority of the sampled video frames in PmLR50 range between 400 and 600 frames. Among a total of five surgical phases, ‘Resecting’ contains the highest number of frames, accounting for 66.61% of the total frames, significantly higher than the 9.38% for ‘Knotting’ and 5.86% for ‘Releasing’ (see the right of Fig. 3). Notably, our dual-rate frame sampling (i.e., 3 fps and 0.33 fps) ensures more than 30 frames of blocking-related operations per surgical procedure, thus providing more balanced samples for various surgical phases.
@@
PM Workflow Recognition Precision $\uparrow$ 88.49 82.78 88.10 84.58 92.36
Recall $\uparrow$ 84.52 70.78 86.32 61.45 86.93
Accuracy $\uparrow$ 92.53 86.59 93.08 84.51 93.83
Jaccard $\uparrow$ 75.12 61.66 79.10 57.98 81.47
@@
Params $\downarrow$ 24.74M 63.12M 24.75M 121.26M 66.62M
Inference Speed (FPS) $\uparrow$ - - - 29.85 35.34
@@
Implementation Details: Our model is trained with the AdamW optimizer for 50 epochs on two NVIDIA 4090 GPUs. The batch size is set to 16, and the initial learning rate is 3e-5. We adopt EfficientNet-B3 (Tan and Le 2021) pre-trained on ImageNet (Deng et al. 2009) as the backbone. The parameters of other components in PmNet are randomly initialized. In our experiments, 20 consecutive frames with frame rate $R=8$ are selected as model inputs in an online manner, with the clip window size $w$ set to 4. During training, we apply color jitter and random horizontal flip for data augmentation. To trade off the accuracy and efficiency of PmNet, we embed four token swap operations in MTE, with the number of SSM blocks set to 2 in CSM.
@@
As illustrated in Table 2, our model shows superior performance in both tasks. Compared to the second-best model Trans-SVNet, our model achieves an improvement of 2.37% Jaccard in PM workflow recognition. Besides, by modeling long-term memory through CSM, PmNet obtains a high accuracy of 98.21% in blocking effectiveness detection. In addition, our model offers certain advantages in terms of parameters and inference efficiency, reaching an inference speed of 35.34 fps for real-time clinical applications. PmNet also outperforms other methods across all phases, which can be attributed to CPS to discriminate between confused intraoperative actions. Fig. 5 visualizes the predictions of our method compared to the latest Surgformer for PM workflow recognition, where PmNet demonstrates better consistency and accuracy across all surgical phases.
@@
Ablations for Each Component. Table 3 ablates the contribution of each component of PmNet to surgical workflow recognition. The experimental results indicate that all components, including MTE, CSM, and CPS play a positive role. The results in the last two rows illustrate that MTE used for capturing short-term temporal information shows the most significant effect, increasing the Jaccard value for PM workflow recognition by 6.44%. Meanwhile, the efficient long-term temporal modeling by CSM and the comparative learning of intraoperative activities by CPS also enhance the validity and robustness of PmNet.
@@ Conclusion
Acknowledgments
The work described in this paper was supported in part by the Research Grants Council of the Hong Kong Special Administrative Region, China, under Project T45-401/22-N; and by the Regional Joint Fund of Guangdong (Guangdong–Hong Kong–Macao Research Team Project) under Grant 2021B1515130003; and by the Shenzhen Science and Technology Program, under Grant RCYX20231211090127030.